An open macroeconomic model for a hypothetical economy is represented as follows

Y= C0 +Io+Go+X0-M, M=mo+m1yd,C=co+c1yd, T=tY and Yd=Y-T

Show that equal change in tax and government expenditure are expansionary to the economy
Derive the equilibrium level of savings in the economy above
Derive the investment multiplier
1
Expert's answer
2015-03-28T10:41:40-0400
The equal change in tax and government expenditure are expansionary for the economy, because if the government expenditure increases (Go to G1), the GDP will increase too, as Y= C0 +Io+Go+X0-M.
So, the equal change in tax and government expenditure will have expansionary effect. In the equilibrium savings are equal to investment, so in our case the equilibrium level of savings is S = Io. Investment multiplier is simply the multiplier effect of an injection of investment into an economy. The investment multiplier in our case will be mi = 1/(1 - c) = 1/(1 - c1).